# Influence of density-dependent diffusion on pattern formation in a bounded habitat

### v3

@@ Abstract
Considering a nonlocal version of the Fisher-KPP equation, we explore the impact of heterogeneous diffusion on pattern formation within a bounded region (refuge). Under homogeneous diffusion, it is established that, nonlocality can lead to spontaneous pattern formation under certain parameter conditions, otherwise, when the homogeneous state is stable, spatial perturbations such as the existence of a refuge (high quality region within a hostile environment) can induce patterns. To understand how density-dependent diffusivity influences these forms of pattern formation, we examine diffusivity reactions to both rarefaction or overcrowding. Additionally, for comparison, we investigate the scenario where diffusivity levels vary spatially, inside and outside the refuge. We find that state-dependent diffusivity affects the shape and stability of patterns, potentially triggering either explosive growth or fragmentation of the population distribution, depending on how diffusion responses to density changes.
@@ I Introduction
A remarkable property of biological systems is the formation of spatial structures. Patterns can emerge by self-organization, as a consequence of the interactions between individuals, without external drivers von2001diversity ; rietkerk2008regular ; rietkerk2002self . In particular, spatially-extended (nonlocal) competitive interactions are sufficient to produce self-organized spatial oscillations murray2002 , whose characteristic wavelength is ruled by the range of these interactions fuentes2003nonlocal ; fuentes2004analytical ; hernandez2004clustering . For instance, plant competition for water, known for generating spatial patterns, can be considered a nonlocal process due to root spatial structure or water diffusive dynamics escaff2015localized ; fernandez2019front . The nonlocality of other elementary processes, such as reproduction and random dispersal, has a less central role but can interfere constructively or destructively with the possibility of pattern formation piva2021interplay .
@@
Although patterns can emerge solely from interactions among individuals, they are naturally affected o even induced by environmental conditions, as these control the rates of the biological processes smith2003animal ; corley1977influence . Such effects can be observed at ecological scales in nature, as well as experimentally in the laboratory. An example is the formation of regular vegetation patterns tarnita2017theoretical in a landscape constituted by patches characterized by different growth rates pimm2019connecting ; maciel2013individual ; wintle2019global . This scenario can also be produced artificially in the lab, as when bacterial colonies are subjected to adverse conditions such as ultra-violet light, except for a protected area (refuge) perry2005experimental . In fact, growth-rate heterogeneity can induce pattern formation, even under conditions that will not give rise to patterns in homogeneous media dornelas2021landscape .
@@
In any case, when addressing pattern formation of biological populations, typically, normal diffusion with constant coefficient is assumed. However, in real environments, not only growth rates, as in the case of a refuge, but also mobility can be heterogeneous, e.g., in active suspensions where movement is influenced by nutrient gradients diluzio2005escherichia , or due to structural features of the environment, as the movement of bacteria in porous media  sosa2017motility ; muskat1937flow . Furthermore, heterogeneous diffusion can also emerge as a reaction arlt2019dynamics that the individuals manifest, for instance, in response to overcrowding or sparsity of a population, favoring, or not, the random motion among other individuals colombo2012nonlinear ; matthysen2005 ; kun2006 ; french2001 ; Cates2010 ; murray2002 ; lopez2006macroscopic ; kenkre2008nonlinearity ; lenzi2001escape ; anteneodo2005non ; anteneodo2007brownian ; newman1980some ; gurtin1977diffusion ; kareiva1983local ; birzu2019genetic . The specific way organisms respond to concentration depends on several conditions and vary from species to species gurtin1977diffusion ; kareiva1983local ; newman1980some ; murray2002 . For instance, in populations of grasshoppers, the diffusion coefficient is enhanced at high densities, where encounters between individuals are more frequent, but in other species, this occurs at low densities murray2002 . Density-dependent factors are also present in migratory dispersal amarasekare2004role ; Matthysen2005density ; Kun2006 .
State-dependent diffusivity in biological population dynamics has been previously studied in the context of critical conditions for survival amarasekare2004role ; colombo2012nonlinear ; maciel2013individual , and also with regard to pattern formation. Let us mention an important study on the distinct roles on pattern formation of Fokker-Planck and Fick’s laws of diffusion, for spatially varying coefficient of diffusion ($D(x)$bengfort2016fokker . Although we will address systems with nonlocality as pattern-formation mechanism, let us mention that there works showing that the effects of nonhomogeneous environments cannot be neglected in systems with Turing instabilities Yizhaq2014 ; Cobbold2015 . Furthermore, in these works the heterogeneities are space-dependent while the density dependencies have not received sufficient attention, as far as we know.
@@
Then, we analyze two classes of heterogeneous diffusivity: state-dependent (where the diffusivity responds to the population density) and spatial-dependent (where diffusivity is associated to the quality of the environment). In both cases there might be a feedback that mitigates or reinforces pattern formation. Moreover, we focus on the effects that diffusive heterogeneities have on the spacial distribution of a population inhabiting a refuge immersed in an adverse environment. We consider as starting point the description of a single-population dynamics given by the spatially-exteded (nonlocal) form of the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) dynamics murray2002 , which includes random movements (normal diffusion) and logistic growth with nonlocal competition. Then we generalize this equation by substituting normal diffusion by each form of heterogeneous diffusion considered. For state-dependent diffusivity, we focus on two functional forms, namely, decay or increase with the population density, reflecting enhanced mobility in response to sparseness and overcrowding, respectively. For this purpose, exponential dependencies are studied as paradigm. For the spatial-dependent case, we consider that diffusivity is associated to the quality of the environment, which is different inside and outside the refuge.
@@ III.1 Survival
The critical size of a refuge, for population survival, is known to depend on growth and diffusion coefficients, both inside and outside the refuge, when discontinuous binary forms of $r(x)$ and $D(x)$ are considered maciel2013individual ; colombo2016intermittent ; colombo2018refuge ; ludwig79 ; skellam91 ; kenkre03 . Succinctly, the critical size is derived assuming small $u$, hence solving Eq. (2) in each domain after discarding nonlinearities, and coupling the solutions through continuity conditions. As the diffusivity is a continuous function in all the studied cases cases, then we impose continuity for $u$ and its derivative at the interfaces. Hence, the critical threshold for survival is approximately
@@ III.2 Linear analysis
which is the growth rate of mode $k$ that can also be obtained by Fourier transforming Eq (7). Since at first order in this approximate approach the diffusivity is nearly constant, let us review the picture known for uniform diffusivity with periodic boundary conditions. If $\lambda(k)>0$, the perturbation grows, with characteristic mode $k^{*}$ which maximizes $\lambda(k)$, providing a characteristic wavelength $\Lambda^{*}=2\pi/k^{*}$. Furthermore, considering the first maximum of Eq. (8), one obtains the critical condition for pattern formation kenkre03 ; colombo2012nonlinear
@@
when $r>0$; otherwise, no oscillations arise. The condition (9) is crucial to produce patterns in a homogeneous environment.
However, in a landscape with heterogeneous growth rate, such as in the case of the refuge with $r(x)$ given by Eq. (1), and constant diffusivity $\bar{D}$, the uniform density state $u_{0}$ can be destabilized even for $\lambda(k)<0$ and damped modes emerge from the complex roots of $\lambda(k)=0$ dornelas2021landscape . In fact, linearization of Eq. (2), at first order in $\varepsilon$ and in $A=r_{in}-r_{out}$, leads to
@@
where $\sim$” indicates Fourier transform and $\psi(x)=A\Theta(|x|-L/2)$. The dominant mode (with lower damping) is given by the root $k=k_{R}\pm ik_{I}$ with the smallest imaginary part (in absolute value). Its imaginary part gives the damping rate while the real part gives the associated wavelength $\Lambda=2\pi/|k_{R}|$. Therefore, depending on the real and imaginary parts of the relevant root, we can observe:
(I) sustained oscillations (when the imaginary part is zero), which corresponds to condition (9);
(II) damped oscillations (when real and imaginary parts are non-null);
(III) absence of oscillations (when the real part is null).
@@ IV Results for state-dependent diffusivity
IV.1 Enhanced reaction to sparsity: $D_{1}(u)=d\exp(-u/\sigma)$
@@
Initially (at $t=0$), the density, hence the diffusivity, are nearly constant (with $D_{c1}<d\exp(u_{0}/\sigma)\simeq 0.036<D_{c2}$). In a first regime ($t\lesssim 1$), the two-level profile of the growth rate $r(x)$ induces essentially a two-level population density, which in turn molds the diffusivity profile. As soon as Eq. (11) holds, steady oscillations emerge, with the characteristic wavelength predicted by the linear analysis, close to $w/0.715$ colombo2012nonlinear .
@@
Approximately, $D_{in}\simeq d\exp(-u_{0}/\sigma)$ in the refuge and $D_{out}\simeq d$ outside far from the interface. Therefore, $\sigma$ affects the internal diffusivity, producing a level that can impact pattern formation, as well as the external population close to the interface, affecting the persistence of the population in this region. For increasing $\sigma$, when the internal diffusivity level exceeds the critical value $D_{c2}\simeq 0.37$, oscillations are spoiled, and the dependence on $u$ is flattened, producing homogeneous diffusivity. In the opposite case of small $\sigma$, the diffusivity is more sensitive to the variations of density and results a density profile with sharp peaks near the interfaces.
@@
Figure 3: Phase diagrams in the plane $\sigma-d$ of $D_{1}(u)$, as in Fig. 9. For the crossed circles a steady state is not observed. Fixed parameters are $r_{in}=1$, $r_{out}=-0.02$.
The dashed and full lines correspond to the conditions given by Eqs. (9) and (11), respectively, using $\bar{D}=D(u_{0})$, which delimit the different phases I-III (see Sec. III). Despite the fluctuations of the density, the approximation of a constant diffusivity $D(u_{0})$ in the internal region gives a prediction in reasonable agreement with the results from simulations. We verified that pattern wavelength agrees with the prediction $\Lambda=2\pi/|k_{R}|$. Nonetheless, in the region where the imaginary part vanishes and sustained oscillations are expected, what actually happens is that a steady state is not attained (crossed symbols). High peaks give rise to a locally low diffusion coefficient, then the fluxes outwards from the peaks are not able to balance the growth of the population and this effect is reinforced the higher the peak, as can be seen in Fig. 4 where a region of the profiles is amplified.
Notice that such feedback does not take place for density-independent diffusivity, in which case a steady state is always achieved. In comparison to the space-dependent case, we also highlight that for $D_{1}(u)$, when stationary oscillations are formed in the refuge, they are strongly damped towards the center. Moreover, the population outside the refuge can attain a moderate level at the interface and decay slowly with the distance to the interface.
Figure 4: Detail of the time evolution if $u(x)$ and $D_{1}(u)$ for $\sigma=0.2$ and $d=1.6$ (corresponding to a crossed circle in Fig 3, and $r_{in}=1$, $r_{out}=-0.02$.
IV.2 Enhanced reaction to overcrowding: $D_{2}(u)=d[1-\exp(-u/\sigma)]$
@@
Although the diffusivity in the outer region becomes lower that in the refuge, oscillations are not formed due to the negative growth rate. Inside the refuge, the diffusivity is nearly constant ($D_{2}\simeq d$), when $\sigma\ll u_{0}$. As $\sigma$ increases, the average level of the diffusivity decreases, favoring oscillations, which tend to become sustained when the diffusivity profile becomes bounded by $D_{c1}$. Furthermore, due to the particular dependency on $u$, the diffusivity is higher at the peaks of $u$ and very small at the valleys, favoring a steady state with barely sustained oscillations, differently to what we observed in the opposite case of reaction to sparsity. Moreover, for sufficiently large $\sigma$, as the amplitude of the sustained oscillation grows, fragmentation of the population inside the refuge can occur (with bumps separated by depopulated regions), a phenomenon observed for the space-dependent profile but not for $D_{1}(u)$.
@@
Figure 6: Phase diagrams in the plane $\sigma-d$ of $D_{2}(u)$, as in Fig. 3. Patterns are formed except in the white region in (a). Fixed parameters $r_{in}=1$, $r_{out}=-0.02$.

### v4

# Influence of density-dependent diffusion on pattern formation in a refuge

@@ Abstract
We investigate a nonlocal generalization of the Fisher-KPP equation, which incorporates logistic growth and diffusion, for a single species population in a viable patch (refuge). In this framework, diffusion plays an homogenizing role, while nonlocal interactions can destabilize the spatially uniform state, leading to the emergence of spontaneous patterns. Notably, even when the uniform state is stable, spatial perturbations, such as the presence of a refuge, can still induce patterns. These phenomena are well known for environments with constant diffusivity. Our goal is to investigate how the formation of winkles in the population distribution is affected when the diffusivity is density-dependent. Then, we explore scenarios in which diffusivity is sensitive to either rarefaction or overcrowding. We find that state-dependent diffusivity affects the shape and stability of the patterns, potentially leading to either explosive growth or fragmentation of the population distribution, depending on how diffusion reacts to changes in density.
@@ I Introduction
A remarkable property of biological systems is the formation of spatial structures. Patterns can emerge by self-organization as a result of specific interactions between individuals, without the need for external drivers von2001diversity ; rietkerk2008regular ; rietkerk2002self . In particular, in the framework of Fisher-type dynamics, which includes logistic growth and diffusion murray2002 , when competitive interactions are spatially extended (nonlocal), they can give rise to self-organized spatial oscillations, with the characteristic wavelength determined by the range of these interactions fuentes2003nonlocal ; fuentes2004analytical ; hernandez2004clustering . For instance, plant competition for water, known for generating spatial patterns, can be considered a nonlocal process due to root spatial structure or water diffusive dynamics escaff2015localized ; martinez2013vegetation . The nonlocality of other elementary processes, such as reproduction and random dispersal, has a less central role but can interfere constructively or destructively with the possibility of pattern formation piva2021interplay . Furthermore, while within Fisher dynamics diffusion can be detrimental to pattern formation, since it promotes homogenization, it can still influence the shape and stability of the resulting patterns, depending on the kind of diffusion (e.g., normal or anomalous) colombo2012nonlinear .
@@
Although patterns can emerge solely from interactions among individuals, they are naturally affected or even induced by environmental conditions, as these control the rates of biological processes smith2003animal ; corley1977influence . Such effects can be observed at ecological scales in nature, as in the case of vegetation patterns induced by spatial heterogeneities sheffer2012 ; tarnita2017 ; pintoramos2022 , as well as artificially produced in the laboratory, such as when bacterial colonies are subjected to adverse conditions such as ultraviolet light, except for a protected area (refuge) perry2005experimental . In fact, growth-rate heterogeneity can induce pattern formation, even under conditions that will not give rise to patterns in homogeneous media dornelas2021landscape .
@@
Furthermore, in real environments, not only growth rates, as in the case of a refuge, but also mobility can be heterogeneous, e.g., in active suspensions where movement is influenced by nutrient gradients diluzio2005escherichia , or due to structural features of the environment, as the movement of bacteria in porous media  sosa2017motility ; muskat1937flow . Furthermore, heterogeneous diffusion can also emerge as a reaction arlt2019dynamics that the individuals manifest, for instance, in response to overcrowding or sparsity of a population, favoring, or not, the random motion among other individuals colombo2012nonlinear ; matthysen2005 ; kun2006 ; french2001 ; Cates2010 ; murray2002 ; lopez2006macroscopic ; kenkre2008nonlinearity ; lenzi2001escape ; anteneodo2005non ; anteneodo2007brownian ; newman1980some ; gurtin1977diffusion ; kareiva1983local ; birzu2019genetic . The specific way organisms respond to concentration depends on several conditions and vary from species to species gurtin1977diffusion ; kareiva1983local ; newman1980some ; murray2002 . For instance, in populations of grasshoppers, the diffusion coefficient is enhanced at high densities, where encounters between individuals are more frequent, but in other species, this occurs at low densities murray2002 . Another source of heterogeneous diffusion, accompanied by bias, is chemotaxis mittal2003motility ; hillen2009user ; arumugam2021keller ; kenkrel2008nonlinearity . Density-dependent factors are also present in migratory dispersal amarasekare2004role ; Matthysen2005density ; kun2006 . State-dependent diffusivity in biological population dynamics has been previously considered in the context of critical conditions for survival amarasekare2004role ; colombo2012nonlinear ; maciel2013individual , and also with regard to pattern formation liu2013 ; liu2016 . Let us mention an important study on the distinct roles on pattern formation of Fokker-Planck and Fick’s laws of diffusion, for spatially varying coefficient of diffusion, $D(x)$ bengfort2016fokker . Although we will address systems with nonlocality as pattern-formation mechanism, let us mention that there are works showing that the effects of nonhomogeneous environments cannot be neglected in systems with Turing instabilities Yizhaq2014 ; Cobbold2015 .
@@
In this work, we analyze two classes of heterogeneous diffusivity: state-dependent (where the diffusivity responds to the population density) and space-dependent (where diffusivity is associated to the quality of the environment). In both cases there might be a feedback that mitigates or reinforces pattern formation. Moreover, we focus on the effects that diffusive heterogeneities have on the spatial distribution of a population inhabiting a refuge immersed in an adverse environment. We consider as starting point the description of a single-population dynamics given by the spatially-exteded (nonlocal) form of the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) dynamics murray2002 , which includes random movements (normal diffusion) and logistic growth with nonlocal competition. Then we generalize this equation by substituting normal diffusion by each form of heterogeneous diffusion considered. For state-dependent diffusivity, we focus on two functional forms, namely, decay or increase with the population density, reflecting enhanced mobility in response to sparseness and overcrowding, respectively. For this purpose, exponential dependencies are studied as paradigm. For the spatial-dependent case, we consider that diffusivity is associated to the quality of the environment, which is different inside and outside the refuge.
@@ III.1 Survival
The critical size of a refuge, for population survival, is known to depend on growth and diffusion coefficients, both inside and outside the refuge, when discontinuous binary forms of $r(x)$ and $D(x)$ are considered maciel2013individual ; colombo2016intermittent ; colombo2018refuge ; ludwig79 ; skellam91 ; kenkre03 . Succinctly, the critical size is derived assuming small $u$, hence solving Eq. (2) in each domain after discarding nonlinearities, and coupling the solutions through continuity conditions. As the diffusivity is a continuous function in all the studied cases cases, then we impose continuity for $u$ and its derivative at the interfaces. Hence, the critical threshold for survival is approximately (see for instance colombo2016intermittent )
@@ III.2 Linear stability analysis
which is the growth rate of mode $k$ that can also be obtained by Fourier transforming Eq (7). Since at first order in this approximate approach the diffusivity is nearly constant, let us review the picture known for uniform diffusivity with periodic boundary conditions. If $\lambda(k)>0$ for some $k>0$, the uniform density state is destabilized by a perturbation and wrinkles develop. The dominant mode with wavelength $\Lambda^{*}=2\pi/k^{*}$, is given by the wavenumber $k^{*}$ which maximizes $\lambda(k)$. For Eq. (8), the first (global) maximum is found at $k^{*}\simeq 1.43\pi/w$, and imposing $\lambda(k^{*})>0$, one obtains the instability condition  kenkre03 ; colombo2012nonlinear
@@
when $r>0$; otherwise, no spatial oscillations arise. Let us highlight that all oscillations observed in the studied system are not temporal but spatial.
The condition (9) is necessary to produce patterns in a homogeneous landscape, with constant $r$. However, in a landscape with heterogeneous growth rate, such as in the case of the refuge with $r(x)$ given by Eq. (1), and constant diffusivity $\bar{D}$, the uniform density state $u_{0}$ can be destabilized even for $\lambda(k)<0$ and (spatially) damped modes emerge from the complex roots of $\lambda(k)=0$ dornelas2021landscape . In fact, linearization of Eq. (2), at first order in $\varepsilon$ and in $A=r_{in}-r_{out}$, leads to
@@
where $\sim$” indicates Fourier transform and $\psi(x)=A\Theta(|x|-L/2)$. To obtain the induced perturbation in real space, we must compute the inverse Fourier transform of Eq. (10). By applying the residue theorem to this integration, it becomes clear that the perturbation takes the form
$\varepsilon(x)=\sum_{j}C_{j}\,e^{(ik_{R,j}-k_{I,j}x},$ (11)
where, the oscillation parameters $k_{R,j}$ (wavenumber) and $k_{I,j}$ (inverse of the decay-length) are the absolute values of the real and imaginary parts of the $j$-th zero of $\lambda(k)$, and the constant coefficients $C_{j}$ depend on the perturbation $\tilde{\Psi}$, which is assumed to be non-periodic (i.e., it does not add any characteristic mode by itself). These roots can be obtained numerically.
The dominant mode (with lower damping) is given by the root $k=k_{R}\pm ik_{I}$ of $\lambda(k)$, with the smallest imaginary part (in absolute value). Its imaginary part gives the damping rate while the real part gives the associated wavelength $\Lambda=2\pi/|k_{R}|$ of the pattern. Therefore, depending on the real and imaginary parts of the relevant root, we can observe:
(I) sustained spatial oscillations (when the imaginary part is zero), which corresponds to condition (9);
(II) damped spatial oscillations (when real and imaginary parts are non-null);
(III) absence of spatial oscillations (when the real part is null).
@@ IV Results for state-dependent diffusivity
IV.1 Enhanced diffusion due to sparsity: $D_{1}(u)=d\exp(-u/\sigma)$
@@
Initially (at $t=0$), the density, hence the diffusivity, are nearly constant, satisfying $D_{c1}<d\exp(-u_{0}/\sigma)\simeq 0.036<D_{c2}$ (according to Eqs. (9)-(12), having used $u_{0}=r=w=1$, $d=1$ and $\sigma=0.3$). In a first regime ($t\lesssim 1$), the two-level profile of the growth rate $r(x)$ induces essentially a two-level population density, which in turn molds the diffusivity profile. As soon as Eq. (12) holds, steady oscillations emerge, with the characteristic wavelength $\Lambda^{*}$ predicted by the linear analysis.
@@
Approximately, $D_{in}\simeq d\exp(-u_{0}/\sigma)$ in the refuge and $D_{out}\simeq d$ outside far from the interface. Therefore, $\sigma$ affects the internal diffusivity, producing a level that can impact pattern formation, as well as the external population close to the interface, affecting the persistence of the population in this region. For increasing $\sigma$, when the internal diffusivity level exceeds the critical value $D_{c2}\simeq 0.37$ (according to Eq. (12), and recalling that we set $r=w=1$), oscillations are spoiled, and the dependence on $u$ is flattened, producing homogeneous diffusivity. In the opposite case of small $\sigma$, the diffusivity is more sensitive to the variations of density and results a density profile with sharp peaks near the interfaces. For even smaller values of $\sigma$ or $d$ than those shown in the figure, the peaks grow without attaining a steady state (not shown).
@@
Figure 3: Phase diagrams in the plane $\sigma-d$ of $D_{1}(u)$. The regions I-III correspond to those described in Sec. III, delimited by dashed and full lines, given by Eqs. (9) and (12), respectively, using $\bar{D}=D_{1}(u_{0})$. The heat plots represent the real $k_{R}$ (a) and imaginary $k_{I}$ (b) parts of the dominant root of $\lambda(k)$. The symbols correspond to the type of density profile observed in the numerical long-time solution of Eq. (2): damped oscillations (filled), uniform (hollow), and for the crossed circles a steady state is not observed. Fixed parameters are $r_{in}=1$, $r_{out}=-0.02$.
The dashed and full lines correspond to the conditions given by Eqs. (9) and (12), respectively, using $\bar{D}=D(u_{0})$, which delimit the different phases I-III (see Sec. III). Despite the oscillations of the density, the approximation of a constant diffusivity $D(u_{0})$ in the internal region gives a prediction in reasonable agreement with the results from simulations. We verified that pattern wavelength agrees with the prediction $\Lambda=2\pi/|k_{R}|$. Nonetheless, in the region where the imaginary part vanishes and sustained oscillations are expected, what actually happens is that a steady state is not attained (crossed symbols). A high peak gives rise to a locally low diffusion coefficient, then the fluxes outwards from the peak are not able to balance the growth of the population and this makes the peak even higher, an effect which is reinforced the higher the peak.
Notice that such feedback does not take place for density-independent diffusivity, in which case a steady state is always achieved. We also highlight that for $D_{1}(u)$, when stationary oscillations are formed in the refuge, they are strongly damped towards the center. Moreover, the population outside the refuge can attain a moderate level at the interface and decay slowly with the distance to the interface.
IV.2 Enhanced diffusion due to overcrowding: $D_{2}(u)=d[1-\exp(-u/\sigma)]$
@@
Although the diffusivity in the outer region becomes lower that in the refuge, oscillations are not formed due to the negative growth rate. Inside the refuge, the diffusivity is nearly constant ($D_{2}\simeq d$), when $\sigma\ll u_{0}$. As $\sigma$ increases, the average level of the diffusivity decreases, favoring oscillations, which tend to become sustained when the diffusivity profile becomes bounded by $D_{c1}$. Furthermore, due to the particular dependency on $u$, the diffusivity is higher at the peaks of $u$ and very small at the valleys, favoring a steady state with barely sustained oscillations, differently to what we observed in the opposite case of reaction to sparsity. Moreover, for sufficiently large $\sigma$, as the amplitude of the sustained oscillation grows, fragmentation of the population inside the refuge can occur (with bumps separated by depopulated regions), a phenomenon observed for the space-dependent profile but not for $D_{1}(u)$. Note that $D_{2}(u)$ tends to zero when $u$ tends to zero, but similar features arise with a baseline different from zero (not shown).
@@
Figure 5: Phase diagrams in the plane $\sigma-d$ of $D_{2}(u)$. The regions I-III correspond to those described in Sec. III, delimited by dashed and full lines, given by Eqs. (9) and (12), respectively, using $\bar{D}=D_{2}(u_{0})$. As in Fig. 3, the heat plots represent the real $k_{R}$ (a) and imaginary $k_{I}$ (b) parts of the dominant root of $\lambda(k)$. The symbols correspond to the type of density profile observed in the numerical long-time solution of Eq. (2): damped oscillations (filled), uniform (hollow). Fixed parameters $r_{in}=1$, $r_{out}=-0.02$.